the standard reduction potentals,E,for the half cells reactions are as


Zn=Zn2++2e;E=+0.76V


Fe=Fe+2+2e;E=0.41V


The emf of the cell reaction


Fe+2+Zn=Fe+Zn is


a)-0.35v


b)+0.35v  d)-1.17v


c)1.17v


 


 


 

Grade:12

1 Answers

askiitian.expert- chandra sekhar
14 years ago

Hi,

Fe+2+Zn=Fe+Zn+2

E cell = -0.41+0.76

           =0.35 V

ans is (b)